OpenStudy (anonymous):

-x^2-7x+9 find the vertex

OpenStudy (amistre64):

the x component of the vertex is the opposite of the middle number divided by 2

OpenStudy (amistre64):

-(-7)/2 = 7/2; now use that to find the value of the function

OpenStudy (amistre64):

-(7/2)^2 -7(7/2) +9 = ?

OpenStudy (amistre64):

i get -27.75

OpenStudy (amistre64):

so vertex = (\(\frac{7}{2},\frac{-111}{4}\)) maybe?

OpenStudy (anonymous):

the correct answer was (7/2,85/4)

OpenStudy (amistre64):

hmm ..... i must miss placed a negative someplace

OpenStudy (anonymous):

here is another one -x^2-7x+7

OpenStudy (amistre64):

-x^2-7x+9; Vx = 7/-2 if i see it right; you sure the correct answer doesnt start with -7/2?

OpenStudy (anonymous):

oops the new problem is -x^2-x+7

OpenStudy (amistre64):

-(x^2 +x -7) = y x^2 +x -7 = -y x^2 +x = -y+7 x^2 +x +(1/2)^2 = -y +7 +(1/2)^2 (x+1/2)^2 = -y +(29/4) -(x+1/2)^2 = y - (29/4) y = -(x+(1/2))^2 + (29/4) vertex = (1/2, 29,4) maybe?

OpenStudy (anonymous):

not correct

OpenStudy (amistre64):

either you got a broken program; or someone broke the math today :)

OpenStudy (anonymous):

correct answer was -1/2,29/4

OpenStudy (amistre64):

thiiiissss close; its the negative in front that is throwing me off ... bummer

OpenStudy (amistre64):

wolfram says my math was right; I just interpreted it wrong afterwards